# Revisiting the Galactic Winds in M82 I: the recent starburst and launch of outflow in simulations

### v2

@@ Abstract
We revisit the launch of the galactic outflow in M82 through hydrodynamic simulations. Employing a sink-particle module, we self-consistently resolve the star formation and feedback processes, avoiding the reliance on various assumed models. We probe the effects of different stellar feedback mechanisms, gas return from star-forming clouds and gas disc mass on the starburst and outflow. Our simulations can generate a starburst that lasts $\sim 25$ Myr, peaking at 20-50 $\rm{M_{\odot}yr^{-1}}$. However, the total stellar mass formed in the starburst often exceeds M82’s estimated value. The outflow’s launch occurs in two stages. Initially, continuous SNe explosions form small bubbles, merging into a super bubble foam composed of warm/hot gas and high-density cool filaments. After $\sim 10$ Myr of SN injection, the super bubble breakout the disc, marking the second stage, which takes $\sim 15$ Myr to develop a kpc-scale outflow. Our simulations reveal that cool filaments within the ISM can survive from the stellar feedback, then were entrained into the outflow and stretched to hundreds pc in length. While the mass loading factor of the well-developed outflow is comparable to M82, the cool gas mass outflow rate is often lower, and its velocity is slower than the estimated value in M82 by $\sim 60\%$. Warm and hot gas are $\sim 25\%$ slower. SN feedback acts as the primary driver of the outflow, while gas return significantly influences the starburst and outflow. Other factors have moderate effects. To address the shortcoming in our results, enhanced SN feedback effect due to clustered SNe is likely necessary.
@@
pubyear: 2015pagerange: Revisiting the Galactic Winds in M82 I: the recent starburst and launch of outflow in simulationsB
@@ 1 Introduction
Since their initial discovery in M82 by Lynds & Sandage (1963), galactic winds and outflows have been observed in both nearby and at high redshift (e.g. McCarthy et al. 1987; Heckman et al. 1990; Martin 1999; Heckman et al. 2000; Rupke et al. 2005; Weiner et al. 2009; Martin et al. 2012; Arribas et al. 2014; Genzel et al. 2014; Heckman et al. 2015; Fluetsch et al. 2019; Keerthi Vasan et al. 2023; Perrotta et al. 2023; Xu et al. 2023a; Carniani et al. 2024; Weldon et al. 2024). Observations spanning multiple bands have revealed that galactic winds are complex, multi-phase structures that extend from a few parsecs to tens of kiloparsecs. These outflows contain a mixture of gas phases, including cold ($\rm{T}\lesssim 10^{3}\,$), cool ($\rm{T}\sim 10^{3}-2\times 10^{4}\,$), warm ($2\times 10^{4}\,\rm{K}\lesssim\rm{T}\lesssim 5\times 10^{5}$), hot ($5\times 10^{5}<\rm{T}<10^{7}\,$), and very hot ($\rm{T}>10^{7}\,$) gaseous, and dust (for more details, see reviews by Veilleux et al. 2005, 2020; Heckman & Thompson 2017; Rupke 2018 and references therein). Observations, analytical models, and simulations have established that stellar feedback, including supernovae, stellar winds, radiation, and possibly cosmic rays, and active galactic nuclei (AGN) feedback can drive galactic-scale outflows, reproducing observed multi-phase structures (e.g., Chevalier & Clegg 1985; Strickland & Stevens 2000; Murray et al. 2005). Extensive research suggests that these outflows play a crucial role in the redistributing of the interstellar medium (ISM), regulating the efficiency of star formation and enriching the circumgalactic medium (CGM) and intergalactic medium (IGM) (Dekel & Silk 1986; Cole et al. 1994; Naab & Ostriker 2017). This work focuses on stellar feedback-driven outflows, widely considered to be the primary mechanism for suppressing star formation in low-mass galaxies.
@@
Despite significant progress over the past six decades, many fundamental questions about galactic winds remain unanswered. Observationally, precisely measuring the key properties of outflows, including velocities, mass, momentum, and energy outflow rates, particularly for warm, cool, and cold phases, remains challenging (Heckman & Thompson 2017; Rupke 2018; Veilleux et al. 2020). Theoretically, while analytical models and simulations can explain hot winds, the mechanisms accelerating cool gas to hundreds of km/s remain uncertain. The roles of radiation pressure and cosmic rays in boosting galactic winds and accelerating warm and cool phases are also open questions (Heckman & Thompson 2017; Zhang 2018). Furthermore, simulations of galactic winds in galaxies with similar stellar and gas mass can yield divergent results due to variations in simulation setups, including SN feedback prescriptions, resolution, and initial ISM conditions.
A deeper understanding of how stellar feedback initiates galactic winds and the underlying processes driving their evolution is essential. Such knowledge will not only help interpret growing observational data but also solidify the modern $\Lambda$CDM paradigm of galaxy formation and evolution (e.g.,Naab & Ostriker 2017). Within this framework, galactic winds are crucial for suppressing star formation in galaxies below the characteristic luminosity $\rm{L_{*}}$, preventing overproduction of faint galaxies (e.g. Dekel & Silk 1986; Cole et al. 1994; Somerville & Davé 2015). Recent cosmological simulations of galaxy formation and evolution, such as Horizon-AGN (Dubois et al. 2014), EAGLE (Schaye et al. 2015), IllustrisTNG (Pillepich et al. 2018), Simba (Davé et al. 2019), have incorporated galactic winds and AGN feedback to successfully reproduce many observed galaxy properties. However, despite addressing similar physical processes, these simulations employ diverse sub-grid models for stellar feedback and galactic winds. EAGLE utilizes stochastic thermal heating from massive stars, Horizon-AGN adopts mechanical feedback (Kaviraj et al. 2017), and IllustrisTNG employs isotropic kinetic winds. Furthermore, the values of key parameters, including kinetic energy fraction, velocity, and mass outflow rate, vary significantly across these codes.
@@
To gain a deeper understanding of galactic winds, a systematic theoretical investigation across multiple scales is necessary. On a parsec scale, detailed studies are required to explore the formation and lifetime of cold gas clumps and filaments in hot winds (e.g. Cooper et al. 2009; McCourt et al. 2015; Scannapieco & Brüggen 2015; Gronke & Oh 2018), thermal instability in multi-phase winds, dust-gas coupling, and the efficiency of radiation pressure and cosmic rays, using both analytical models and simulations. Concurrently, global models and simulations spanning kpc scales are needed to examine the combined effects of these processes in realistic galaxy environments. Nearby galaxies with extensive observations, such as M82, provide ideal laboratories for testing these global simulations.
@@
Numerous simulations of M82-like galaxies have been conducted over the past four decades. Early studies often employed one- and two-dimensional, axisymmetric simulations to investigate the development of galactic winds driven by nuclear starbursts in isothermal interstellar medium (e.g. Tomisaka & Ikeuchi 1988; Suchkov et al. 1994, 1996; Tenorio-Tagle & Muñoz-Tuñón 1998; Mac Low & Ferrara 1999; Strickland & Stevens 2000; Tenorio-Tagle et al. 2003; Strickland & Heckman 2009). These idealized simulations have validated the Chevalier & Clegg (1985) model in the adiabatic phase and demonstrated that the properties of the wind, such as dynamics, morphology, and X-ray emission, are influenced by factors such as the distribution of the interstellar medium, the ambient gaseous halo, and the distribution and clustering of young stars.
@@
More recently, three-dimensional simulations (e.g. Cooper et al. 2008; Rodríguez-González et al. 2008; Melioli et al. 2013) have been employed to model M82’s winds. Cooper et al. (2008) demonstrated that H-$\alpha$ filaments form from disk clouds entrained and stretched by the hot wind, emphasizing the role of the inhomogeneous disk in shaping the wind’s filamentary structure. Rodríguez-González et al. (2008) and Melioli et al. (2013) highlighted the importance of radiative interactions between winds from super star clusters (SSCs), as well as the number, distribution, and buildup of SSCs, in developing H-$\alpha$ filaments. Very recently, Schneider et al. (2020) and Schneider & Mao (2024) conducted three-dimensional simulations of outflows driven by clustered supernova feedback in a large domain ($\sim 20$ kpc) with high resolution ($\sim 5$ pc). These simulations successfully produced multiphase outflows that closely resemble M82’s outflow in many aspects..
@@
However, most previous simulations of M82’s winds have some limitations due to oversimplified assumptions and factors. A critical issue is the unresolved nature of the recent starburst in the central disk region. Instead of self-consistently resolving the starburst, previous simulations often manually introduce star clusters with varying total masses (1 to 6 ×$\times 10^{8}\rm{M_{\odot}}$) in different studies Cooper et al. 2008; Rodríguez-González et al. 2008; Melioli et al. 2013; Schneider et al. 2020. spatial distributions, and age distributions. Additionally, assumptions about M82’s properties, such as its mass model and gas distribution, are often outdated. Recent observations have revealed discrepancies with earlier studies, particularly regarding the rotation curve ( Greco et al. 2012 and gas distribution (e.g., Leroy et al. 2015), etc. Furthermore, the resolution and numerical methods employed in many previous simulations, including stellar evolution, radiative cooling, radiative pressure, and heating, can be further refined.
@@
To enhance our understanding of galactic winds, we revisit M82’s outflows using high-resolution three-dimensional hydrodynamic simulations that incorporate updated knowledge of M82, a more realistic representation of the recent nuclear starburst, and refined numerical treatments of relevant physical processes. This work, the first in a series, focuses on resolving the nuclear starburst and the subsequent launch of kpc-scale galactic outflows. Our second work (Li et al. 2024) delves into the development of multiphase outflows and the impact of various factors on their properties. This paper is organized as follows. Section 2 outlines our methodology. Section 3 presents the properties of the starburst in our simulations. The launch of kpc-scale multiphase outflows is described in Section 4. Section 5 discusses our findings, including comparisons with previous work and limitations. Finally, Section 6 summarizes our findings.
@@ 2.1 Mass model of M82
Recent observational advances have provided more detailed estimates of the mass and distribution of baryonic matter, including stars and various gas phases. Furthermore, recent observations of the rotation curve suggest the presence of a dark matter halo, which was not considered in earlier studies. We construct a mass model for M82 that incorporates these latest observations. While significant uncertainties remain in the masses of gas and stars, we account for these by varying parameters in our models. Our mass models generally follow the approach of previous studies (e.g. Strickland & Stevens 2000; Cooper et al. 2008; Melioli et al. 2013; Schneider et al. 2020). Our mass models mainly account for the contribution of a stellar disk, a gas disk, and a dark matter halo. The latter is not considered in earlier works such as Strickland & Stevens (2000); Cooper et al. (2008); Melioli et al. (2013), but is needed to reproduce the latest observed rotation curve of M82 in Greco et al. (2012).
We adopt a Miyamoto & Nagai 1975 profile profile for the stellar disk, with a mass of, with a mass of $3.3\times 10^{9}\,\rm{M_{\odot}}$, a scale length of 1200 pc, and a scale height of 200 pc. M82 also consists of a bulge with a mass of $\sim 5.0\times 10^{8}\rm{M_{\odot}}$ in the central 1 kpc region. A considerable fraction of stars in the bulge is likely formed in the recent star bursts, considering the spatial overlapping. Since this work focuses on simulating the recent starburst, we incorporate the bulge mass into the gas component, that is, the gas disk 2 in Table 1. The dark matter halo in our model follows a core-NFW profile with a mass of $\rm{M_{dmh}}\approx 6.0\times 10^{10}\,M_{\odot}$. The gas disk is modeled using a triple Miyamoto-Nagai profile similar to (Smith et al. 2015), with a mass of $\rm{M_{gd}}\approx 2.5,3.0,3.5\times 10^{9}\,\rm{M_{\odot}}$. We manage to determine the related parameters for each component so that the total rotation curve predicted by these components can reproduce the result in Greco et al. (2012), as Figure 1 shows. Detailed information about the mass model is provided in Appendix A.
@@ 2.2 Initial conditions in simulation
The density profile of the gas disc component in the state of hydrostatic equilibrium under the galaxy gravity potential is given by
$\begin{split}\rho(r,z)=&\rho_{0}\ \times\\ &\rm{exp}\left[-\frac{\Phi_{tot}(r,z)-e^{2}\Phi_{tot}(r,0)-(1-e^{2})\Phi_{tot}(0,0)}{c_{s}^{2}}\right],\end{split}$ (1)
@@
$e=e_{rot}\,\rm{exp}(-z/z_{rot})$ (2)
where $e_{rot}$ in the above equation is chosen to be 0.92 for a thick disk. For the scale height $\rm{z}_{rot}$, we choose a value of $\rm{z}_{rot}=5\,\rm{kpc}$, following the literature (Suchkov et al. 1994; Strickland & Stevens 2000; Cooper et al. 2008). In addition to the gas disc, we also include a hot gas halo that follows a profile similar to that of Eqn. 1 but with a much smaller $e_{rot}$ , which is consistent with previous studies(Strickland & Stevens 2000; Cooper et al. 2008). The parameters for the isothermal disk and the hot gas halo are listed in Table 1. Note that $c_{s}$ stands for the effective sound speed derived from the combined energy density of both the dynamic pressure provided by turbulence and the static pressure provided by internal energy, since the temperature of the gas disk quickly drops below $10^{4}$ K where radiative cooling applies.
Table 1: The initial central density, $\rho_{0}$, rotation factor, $e_{rot}$, and effective sound speed of the gas disc and halo. Gas disc 1 represents the dominate gas disc component. Gas disc 2 actually is used to mimic the potential of bulge, see text for more details.
Component $\rho_{0}$ $e_{rot}$ $c_{s}$
@@
gas disk 1 190 0.92 $3.6\times 10^{6}$
gas disk 2 70 0.5 $5.7\times 10^{6}$
@@
Observations reveal a turbulent interstellar medium (ISM) in M82, with non-thermal velocity dispersions of $\sim 20\,\rm{km\ s^{-1}}$ in molecular clouds (Westmoquette et al. 2012). Turbulence likely plays a significant role in both star formation and disk support. To account for this, we introduce turbulence into the initial gas disk. Unlike Cooper et al. (2008), who employed a static, inhomogeneous ISM with a lognormal probability density function, we perturb the velocity field using an energy spectrum function, given by
@@
$(\delta\textbf{u}(r))^{2}\right>\propto r^{p-1}.$ (4)
For typical Kolmogorov turbulence, $p=\frac{5}{3}$. However, turbulence in the ISM is generally supersonic (e.g.Elmegreen & Scalo 2004; Federrath et al. 2010a), with $p$ closer to 2 (e.g.Boldyrev 2002; Federrath 2013). Given the mean velocity dispersion reported by Westmoquette et al. (2012), the ISM in the star-forming region of M82 is in a supersonic turbulent state. We initialize the gas velocity field as a superposition of rotational and turbulent components. This gas distribution is then evolved without cooling or feedback for 2.5 Myr to generate an inhomogeneous disk, which serves as the initial condition for our simulations.

### v4

@@ Abstract
We revisit the launch of the galactic outflow in M82 using hydrodynamic simulations. Employing a sink-particle module, we self-consistently resolve star formation and feedback, avoiding reliance on simplified models. We investigate the effects of stellar feedback mechanisms, gas return from star-forming clouds, and disk mass on the starburst and outflow. Our simulations generate a starburst lasting $\sim 25$ Myr, peaking at 20–50 $\rm{M_{\odot},yr^{-1}}$, although the total stellar mass often exceeds M82’s estimated value. The outflow develops in two stages: initially, continuous SNe form small bubbles that merge into a superbubble containing warm/hot gas and intermediate- to high-density cool filaments. After $\sim 10$ Myr, the superbubble breaks out of the disk, and within $\sim 15$ Myr a kpc-scale outflow forms. Cool filaments survive stellar feedback, become entrained in the wind, and stretch to hundreds of parsecs. Transport from the cool ISM is the dominant net contributor to the total mass of the cool phase in the outflow, whereas transfers from hotter phases, such as through condensation or precipitation, provide only a minor net contribution, likely offset by simultaneous transfer from the cool phase back to hotter phases. While the mass loading factor is comparable to M82, the cool gas outflow rate and velocity are lower, with velocities $\sim 60\%$ below observed values; warm and hot gas are $\sim 25\%$ slower. SN feedback is the primary driver, and gas return significantly influences the starburst and outflow, while other factors are secondary. Stronger clustered SN feedback is likely required to better match observations.
@@
pubyear: 2015pagerange: Revisiting the Galactic Winds in M82 I: the recent starburst and launch of outflow in simulationsC
@@ 1 Introduction
Since the initial discovery in M82 by Lynds and Sandage (1963), galactic winds and outflows have been observed in both nearby (e.g., McCarthy et al. 1987; Heckman et al. 1990; Martin 1999; Heckman et al. 2000; Rupke et al. 2005; Arribas et al. 2014; Heckman et al. 2015; Fluetsch et al. 2019; Xu et al. 2023b) and high redshift universe(e.g. Weiner et al. 2009; Martin et al. 2012; Genzel et al. 2014; Keerthi Vasan et al. 2023; Perrotta et al. 2023; Carniani et al. 2024; Weldon et al. 2024). Observations spanning multiple bands have revealed that galactic winds are complex, multi-phase structures that extend from a few parsecs to tens of kiloparsecs. These outflows contain a mixture of gas phases, including cold ($\rm{T}\lesssim 10^{3}\,$K), cool ($\rm{T}\sim 10^{3}-2\times 10^{4}\,$K), warm ($2\times 10^{4}\,\rm{K}\leq\rm{T}<5\times 10^{5}$K), hot ($5\times 10^{5}\leq\rm{T}<10^{7}\,$K), and very hot ($\rm{T}\geq 10^{7}\,$K) gaseous, and dust (for more details, see reviews by Veilleux et al. 2005, 2020; Heckman and Thompson 2017; Rupke 2018 and references therein). Observations, analytical models, and simulations have established that stellar feedback, including supernovae, stellar winds, radiation, and possibly cosmic rays, and active galactic nuclei (AGN) feedback can drive galactic-scale outflows, reproducing observed multi-phase structures (e.g., Chevalier and Clegg 1985; Strickland and Stevens 2000; Murray et al. 2005). Extensive research suggests that these outflows play a crucial role in the redistribution of the interstellar medium (ISM), regulating the efficiency of star formation and enriching the circumgalactic medium (CGM) and intergalactic medium (IGM) (Dekel and Silk 1986; Cole et al. 1994; Naab and Ostriker 2017). This work focuses on stellar feedback driven outflows, widely considered to be the primary mechanism for suppressing star formation in low-mass galaxies.
@@
Despite significant progress over the past six decades, many fundamental questions about starburst driven galactic winds remain unanswered. Observationally, precisely measuring the key properties of outflows, including velocities, mass, momentum, and energy outflow rates, particularly for warm, cool, and cold phases, remains challenging (Heckman and Thompson 2017; Rupke 2018; Veilleux et al. 2020). Theoretically, while analytical models and simulations can explain hot winds, the origin of cool gas and the mechanisms accelerating cool gas to hundreds of km/s remain uncertain (e.g. Gronke and Oh 2018; Fielding and Bryan 2022). The roles of radiation pressure and cosmic rays in boosting galactic winds and accelerating warm and cool phases are also open questions (Heckman and Thompson 2017; Zhang 2018). Furthermore, simulations of galactic winds in galaxies with similar stellar and gas mass can yield divergent results due to variations in simulation setups, including SN feedback prescriptions, resolution, and initial ISM conditions.
[…]
@@
To gain a deeper understanding of galactic winds, a systematic theoretical investigation across multiple scales is necessary. On a parsec scale, detailed studies are required to explore the formation and lifetime of cold gas clumps and filaments in hot winds (e.g. Cooper et al. 2009a; McCourt et al. 2015; Scannapieco and Brüggen 2015; Gronke and Oh 2018), thermal instability in multi-phase winds, dust-gas coupling, and the efficiency of radiation pressure and cosmic rays, using both analytical models and simulations. Concurrently, global models and simulations spanning kpc scales are needed to examine the combined effects of these processes in realistic galaxy environments. Nearby galaxies with extensive observations, such as M82, provide ideal laboratories for testing these global simulations.
@@
Numerous simulations of M82-like galaxies have been conducted over the past four decades. Early studies often employed one- and two-dimensional, axisymmetric simulations to investigate the development of galactic winds driven by nuclear starbursts in isothermal interstellar medium (e.g. Tomisaka and Ikeuchi 1988; Suchkov et al. 1994, 1996; Tenorio-Tagle and Muñoz-Tuñón 1998; Mac Low and Ferrara 1999; Strickland and Stevens 2000; Tenorio-Tagle et al. 2003; Strickland and Heckman 2009). These idealized simulations have validated the Chevalier and Clegg (1985) model in the adiabatic phase and demonstrated that the properties of the wind, such as dynamics, morphology, and X-ray emission, are influenced by factors such as the distribution of the interstellar medium, the ambient gaseous halo, and the distribution and clustering of young stars.
@@
More recently, three-dimensional simulations (e.g. Cooper et al. 2008; Rodríguez-González et al. 2008; Melioli et al. 2013) have been employed to model the galactic wind in M82-like galaxies. Cooper et al. (2008) suggested that H-$\alpha$ filaments form from disk clouds entrained and stretched by the hot wind, emphasizing the role of the inhomogeneous disk in shaping the wind’s filamentary structure. Rodríguez-González et al. (2008) and Melioli et al. (2013) highlighted the importance of radiative interactions between winds from super star clusters (SSCs), as well as the number, distribution, and buildup of SSCs, in developing H-$\alpha$ filaments. Very recently, Schneider et al. (2020) and Schneider and Mao (2024) conducted three-dimensional simulations of outflows driven by clustered supernova feedback in a large domain ($\sim 20$ kpc) with high resolution ($\sim 5$ pc). These simulations successfully produced multiphase outflows that closely resemble M82’s outflow in many aspects.
@@
However, as most prior simulations of galactic winds in M82-like galaxies focus on how different physical mechanisms influence the overall properties of outflows, they often adopt some simplified assumptions and models, which limit their ability to compare directly to M82. A common factor is that the recent starburst in the central disk region is not resolved self-consistently. Instead, previous simulations often manually introduce star clusters with varying total masses (1 $\times 10^{8}$ to 6 $\times 10^{8}\rm{M_{\odot}}$), spatial distributions, and age distributions in different studies (Cooper et al. 2008; Rodríguez-González et al. 2008; Melioli et al. 2013; Schneider et al. 2020). Additionally, assumptions about the M82-galaxy properties, such as its mass model and gas distribution, are often outdated from recent observations of M82, which have revealed discrepancies with earlier studies, particularly regarding the rotation curve ( Greco et al. 2012), dark matter halo and gas distribution (e.g., Leroy et al. 2015). Furthermore, the resolution and numerical methods employed in many previous simulations, including stellar evolution, radiative cooling, radiative pressure, and heating, can be further refined.
@@
To enhance our understanding of galactic winds, we revisit M82’s outflows using high-resolution three-dimensional hydrodynamic simulations that incorporate updated knowledge of M82, a more realistic representation of the recent nuclear starburst, and refined numerical treatments of relevant physical processes. This work, the first in a series, focuses on resolving the nuclear starburst and the subsequent launch of kpc-scale galactic outflows. Our second work (Li et al. 2025) delves into the development of multiphase outflows and the impact of various factors on their properties. This paper is organized as follows. Section 2 outlines our methodology. Section 3 presents the properties of the starburst in our simulations. The launch of kpc-scale multiphase outflows is described in Section 4. Section 5 discusses our findings, including comparisons with previous work and limitations. Finally, Section 6 summarizes our findings.
@@ 2.1 Mass model of M82
Recent observational advances have provided more detailed estimates of the mass and distribution of baryonic matter, including stars and various gas phases. Furthermore, recent observations of the rotation curve suggest the presence of a dark matter halo, which was not considered in earlier studies. We construct a mass model for M82 that incorporates these latest observations. While significant uncertainties remain in the masses of gas and stars, we account for these by varying parameters in our models. Our mass models generally follow the approach of previous studies (e.g. Strickland and Stevens 2000; Cooper et al. 2008; Melioli et al. 2013; Schneider et al. 2020). Our mass models mainly account for the contribution of a stellar disk, a gas disk, and a dark matter halo. The latter is not considered in earlier works such as Strickland and Stevens (2000); Cooper et al. (2008); Melioli et al. (2013), but is needed to reproduce the latest observed rotation curve of M82 in Greco et al. (2012).
We adopt a Miyamoto and Nagai 1975 profile for the stellar disk, with a mass of $3.3\times 10^{9}\,\rm{M_{\odot}}$, a scale length of 1200 pc, and a scale height of 200 pc. M82 also consists of a bulge with a mass of $\sim 5.0\times 10^{8}\rm{M_{\odot}}$ in the central 1 kpc region. A considerable fraction of stars in the bulge is likely formed in the recent star bursts, considering the spatial overlapping, and since this work focuses on simulating the recent starburst, we incorporate the bulge mass into the gas component, that is, the gas disk 2 in Table 1. The gas disk is modeled using a triple Miyamoto-Nagai profile similar to (Smith et al. 2015), with a mass of $\rm{M_{gd}}\approx 2.5,3.0,3.5\times 10^{9}\rm{M_{\odot}}$, and a half-mass radius of 1 kpc and a ellipticity of 0.3. The dark matter halo in our model follows a core-NFW profile with a mass of $\rm{M_{dmh}}\approx 6.0\times 10^{10}\,M_{\odot}$. We manage to determine the related parameters for each component so that the total rotation curve predicted by these components can reproduce the result in Greco et al. (2012), as Figure 1 shows. Detailed information about the mass model is provided in Appendix A.
@@ 2.2 Initial conditions in simulation
The density profile of the gas disk component in the state of hydrostatic equilibrium under the galaxy gravity potential is given by
$\begin{split}\rho(r,z)=&\rho_{0}\ \times\\ &\rm{exp}\left[-\frac{\Phi_{tot}(r,z)-e^{2}\Phi_{tot}(r,0)-(1-e^{2})\Phi_{tot}(0,0)}{c_{\rm{{s,eff}}}^{2}}\right],\end{split}$ (1)
@@
$e=e_{\rm{rot}}\,\rm{exp}(-z/z_{rot})$ (2)
where $e_{\rm{rot}}$ in the above equation is chosen to be 0.92 for a thick disk. For the scale height $\rm{z}_{rot}$, we choose a value of $\rm{z}_{rot}=5\,\rm{kpc}$, following the literature (Suchkov et al. 1994; Strickland and Stevens 2000; Cooper et al. 2008). In addition to the gas disc, we also include a hot gas halo that follows a profile similar to that of Eqn. 1 but with a much smaller $e_{\rm{rot}}$ , which is consistent with previous studies(Strickland and Stevens 2000; Cooper et al. 2008). The parameters for the isothermal disk and the hot gas halo are listed in Table 1. Note that $c_{\rm{s,eff}}$ stands for the effective sound speed that correspond to the combined energy density of both the non-thermal pressure provided by turbulence and the static pressure provided by internal energy.
Table 1: The initial central density, $\rho_{0}$, rotation factor, $e_{\rm{rot}}$, and effective sound speed of the gas disc and halo. Gas disc 1 represents the dominate gas disc component. Gas disc 2 actually is used to mimic the potential of bulge, see text for more details.
Component $\rho_{0}$ $e_{\rm{rot}}$ $c_{\rm{s,eff}}$
@@
gas disk 1 (extended disk) 190 0.92 $3.6\times 10^{6}$
gas disk 2 (bulge) 70 0.5 $5.7\times 10^{6}$
@@
Observations reveal a turbulent interstellar medium (ISM) in M82, with non-thermal velocity dispersions of $\sigma\sim 20\,\rm{km\ s^{-1}}$ in molecular clouds (Westmoquette et al. 2012). Turbulence likely plays a significant role in both star formation and disk support. To account for this, we introduce turbulence into the initial gas disk. Unlike Cooper et al. (2008), who employed a static, inhomogeneous ISM with a lognormal probability density function, we perturb the velocity field using an energy spectrum function, given by
@@
$\left<(\delta u(r))^{2}\right>\propto r^{p-1}.$ (4)
For typical Kolmogorov turbulence, $p=\frac{5}{3}$. However, turbulence in the ISM is generally supersonic (e.g.Elmegreen and Scalo 2004; Federrath et al. 2010a), with $p$ closer to 2 (e.g.Boldyrev 2002; Federrath 2013). Given the mean velocity dispersion reported by Westmoquette et al. (2012), the ISM in the star-forming region of M82 is in a supersonic turbulent state. We initialize the velocity of each grid cell as the superposition of rotational and turbulent components, and the latter is assumed to have a velocity energy spectrum with $p=2$, with a velocity dispersion magnitude $\sigma=20\,\rm{km\ s^{-1}}$. The velocity field of the turbulence adopted in our simulation is calculated using the Fourier transform. It is then applied only to the gas disk component.
Then, to set up a stable disk, the static pressure of the gas disc is determined by subtracting the non-thermal pressure provided by turbulence from the effective pressure that calculated using the effective sound speed. The non-thermal pressure provided by turbulence takes the form of
$p_{nt}(v_{tur})=\frac{1}{2}\rho{v_{tur}}^{2}$ (5)
while the pressure of ideal gas takes the form of
$p(c_{s})=\frac{1}{\gamma}\rho c_{s}^{2}$ (6)
The static pressure of gas in disc is then given by
$p_{\rm{static}}=p(c_{\rm{s\_eff}})-p_{nt}(v_{tur})$ (7)
For gas in the hot halo, turbulence is absent, and the static pressure is determined by the effective sound speed. At each grid point, the total gas density is given by the sum of the densities of the three components listed in Table 1, while the static pressure is obtained by summing the static pressures of these components. The gas distribution is then evolved without cooling or feedback for 2.5 Myr to generate an inhomogeneous disk, which serves as the initial condition for our simulations.